3. Calculate the change of entropy for surrounding for the following reaction at room temperature and 1atm.

CH4(g)+2 O2(g)→CO2(g)+2 H2O(l) ΔH°=−890.3 kJ
1
Expert's answer
2020-04-13T02:25:09-0400

∆G=∆H/T=-890.3/(25+273)=2.98758389